OpenStudy (anonymous):

What is the solution of the system of equations? -3x-4y-3z=-7 2x-6y+2z=3 5x-2y+5z=9

OpenStudy (anonymous):

Multiply our first equation \(-3x-4y-3z=-7\) by \(-1\) to yield \(3x+4y+3z=7\). Combine with our second equation \(2x-6y+2z=2\) to yield \(5x-2y+5z=9\). Notice something interesting? That's identical to our third equation, which means the system has infinitely many solutions.

OpenStudy (anonymous):

i got one solution

OpenStudy (anonymous):

i think you copied one equation wrong @oldrin.bataku, the second eqn is =3, not =2

OpenStudy (anonymous):

i get x=9/26, y=-801/338, z=-63/338

OpenStudy (anonymous):

@ktklown good catch. If you combine with \(2x-6y+2z=3\) you actually end up with \(5x-2y+5z=10\). Since we know that \(5x-2y+5z=5x-2y+5z\), we may substitute our equations to yield \(9=10\) -- which is clearly false. There are no solutions.